OpenStudy (anonymous):

A copy machine makes 165 copies in 3 minutes and45 seconds. How many copies does it make per minute?

OpenStudy (anonymous):

hint time t=3 minutes and45 seconds t=3 min x 60 sec/min +45 sec t=180+45=225 sec t= 225 secx(1min/60sec)=3.75 min now that there is 165 copies How many copies does it make per minute?

OpenStudy (anonymous):

@krissy7606

OpenStudy (anonymous):

alternate solution change 45 sec to min then add it to 3 min... t=3 min + 45(1/60) = 3.75 min now that there is 165 copies How many copies does it make per minute?